OpenStudy (tia_sweetheart17):

graph the linear equation. x-2y=-8

OpenStudy (ranga):

One way to quickly graph it is to find the x and y intercepts. To find x-intercept, set y = 0 and solve for x. To find y-intercept, set x = 0 and solve for y. Put a point for the x-intercept on the x-axis. Put a point for the y-intercept on the y-axis. Use a ruler and draw a line passing through those two points.

OpenStudy (ranga):

x-2y=-8 To find x-intercept, set y = 0 and solve for x. x - 0 = -8 x intercept = -8 To find y-intercept, set x = 0 and solve for y. 0 - 2y = -8 y = 4 y intercept = 4 Draw the line.